OpenStudy (anonymous):

proof LIM nto∞ 0.99...9=1?

OpenStudy (anonymous):

are you asking why .99999999999999999999... = 1?

OpenStudy (anonymous):

let .99999999999... =x 10x= 9.99999999999999... x=9.999999999999999999... 10x-x=9 9x=9 x=1 1=.999999999999999999999...

OpenStudy (anonymous):

\[\lim_{n \rightarrow ?\infty}n \sqrt{n ^{2}-n} =\frac{ 1 }{ 2 }\]

hartnn (hartnn):

put y=1/n when n->infinity, y->0 replace all n by 1/y

OpenStudy (experimentx):

there are lot's of reasons ... watch this out http://www.youtube.com/watch?v=TINfzxSnnIE

OpenStudy (experimentx):

and check this out http://math.stackexchange.com/questions/11/does-99999-1